# Attention (as Discrete-Time Markov) Chains

### v1

@@ Abstract
We introduce a new interpretation of the attention matrix as a discrete-time Markov chain. Our interpretation sheds light on common operations involving attention scores such as selection, summation, and averaging in a unified framework. It further extends them by considering indirect attention, propagated through the Markov chain, as opposed to previous studies that only model immediate effects. Our main observation is that tokens corresponding to semantically similar regions form a set of metastable states, where the attention clusters, while noisy attention scores tend to disperse. Metastable states and their prevalence can be easily computed through simple matrix multiplication and eigenanalysis, respectively. Using these lightweight tools, we demonstrate state-of-the-art zero-shot segmentation. Lastly, we define TokenRank—the steady state vector of the Markov chain, which measures global token importance. We demonstrate that using it brings improvements in unconditional image generation. We believe our framework offers a fresh view of how tokens are being attended in modern visual transformers.
@@ 1 Introduction
We introduce the interpretation of individual attention matrices as discrete-time Markov chains (DTMC) (Figure 1), where the (non-negative) attention weights indicate transition probabilities between states that correspond to tokens. This framework, dubbed as Attention Chains, allows analyzing token importance through multiple chain transitions, even though the attention operation models only direct relationships. This is analogous to Google’s seminal paper on DTMC, PageRank (Page et al., 1999), stating that simple counting of incoming hyperlinks does not generally correspond to page importance. Similarly, we argue that considering direct token relationships is not necessarily best for evaluating importance.
@@
Based on this insight, we propose a simple yet powerful operation—considering several transitions, or bounces, of the Markov Chain. This consolidates metastable states and helps filter out noise in the attention signal. As we demonstrate, using multiple bounces yields state-of-the-art results for zero-shot segmentation on a commonly used benchmark. Metastable states can also be identified via eigenanalysis. Specifically, we demonstrate that attention heads representing chains with higher 2nd largest eigenvalues can be more effectively leveraged for downstream tasks, as we show for image segmentation.
@@ 2 Related Work
Due the importance of the attention operation in transformers, studies have proposed modeling them using various mathematical frameworks. For example, Ramsauer et al. (2020) relate attention to classic Hopfield networks, Tsai et al. (2019) frame attention as a kernel smoother, Schlag et al. (2021) find its relation to classic fast weight controllers, or Raghu et al. (2021) compare vision transformers to CNNs. El et al. (2025) discuss the mathematical equivalence between message passing in graph neural networks and the self-attention operation in transformers. Our work proposes to interpret the attention mechanism as a DTMC and we demonstrate the usefulness of this approach in various applications.
@@ 3.1 Discrete-Time Markov Chains (DTMC)
where $\mathbf{v}^{T}_{n=0}$ is an initial state summing up to one. The computation is usually terminated once the norm falls below a specific threshold $||\mathbf{v}^{T}_{n+1}-\mathbf{v}^{T}_{n}||^{2}_{2}<\tau$ or after some fixed number of iterations.
@@ 3.2 PageRank Algorithm
where $10<\alpha<$ 1 is a hyper parameter controlling the transitioning probability of “teleportation” into a random state, and $\mathbf{e}$ is a column vector of all ones. Applying the power method on $\mathbf{P}^{\prime}$ results in a unique vector $\vec{\mathbf{\pi}}$, the global PageRank vector, inducing the order $i\preceq_{\vec{\mathbf{\pi}}}j\quad\Longleftrightarrow\quad\vec{\mathbf{\pi}}_{i}\leq\vec{\mathbf{\pi}}_{j}$.
Figure 2: Illustration of higher order effects. Left: Attention matrix $A$ with sequence length 5. Middle: A DTMC with transition probabilities defined by matrix $A$, where only strong connections are shown. Right (One-Hot): To evaluate where state-4 attends to, we can iterate using the power method once starting from a one-hot vector ($n=0$), which results in the row-select operation ($n=1$). However, this first-order approximation is insufficient since state-0 mostly transitions to state-3 and, therefore, state-4 indirectly attends state-3. This becomes evident as we iterate further ($n=2$). Right (Uniform): To compute a global token ranking, we can iterate starting from a uniform state ($n=0$), resulting in a per-column sum operation ($n=1$). This indicates state-0 as most important because many states have a high probability of transitioning into state-0. However, state-0 maps to state-3 with high probability, and state-3 maps to state-4 with high probability. Therefore, the importance of state-4 should be elevated. When considering the second bounce ($n=2$), more probability mass is directed into state-3, and with a sufficient number of iterations the steady state ($v^{T}_{ss}$) ranks state-4 as the most important state globally, which aligns with the intuition above.
@@ 4 Attention as a Discrete-Time Markov Chain
where $h$ is the number of heads, $s_{1,2}$ are the token sequence lengths, and $d_{h}$ is a sub-space of the full embedding dimension $d$. Consequently, we interpret the attention matrix as a DTMC, going beyond the conventional attention mechanism in transformers, which only models direct interaction between queries and keys. Our discussion is limited to attention blocks of equal sequence size (e.g. self-attention, or hybrid attention blocks), unless otherwise stated. We start by formalizing some common operations performed on attention matrices using the DTMC formalism (Section 4.1). While the previously proposed attention operations can be formulated with a single bounce of the Markov chain, we go beyond this and iterate through the Markov chain. Specifically, we describe three new operations enabled by this framework: multi-bounce attention (Section 4.2); where higher-order attention effects are taken into account, TokenRank (Section 4.3); the token-space equivalent of PageRank, and $\lambda_{2}$ weighting; an improved attention head weighting scheme that is based on the second largest eigenvalue of the chain.
@@ 4.1 Interpretation of Existing Attention Operations
Thus, our formulation allows us to reinterpret the selection operation as the first bounce transition from a one-hot vector, a view that we will further motivate in Section 4.2.
Summing each column of $\mathbf{A}$ results in a row vector $\mathbf{v_{s}}^{T}$ that aggregates attention given to each token from all other tokens. In our formulation, $\mathbf{v_{s}}^{T}$ can be obtained by performing a single state transition: $\mathbf{v}^{T}_{s}=\frac{1}{n}\mathbf{e}^{T}\mathbf{A}$, where $\mathbf{e}^{T}$ is a row vector of ones. This operation can be a useful first-order approximation in ranking tokens’ importance globally, as it equates to starting from an equilibrium state and transitioning through the DTMC process once. However, we show this straight-forward approach under-performs compared to the global TokenRank vector (Section 4.3) that takes into account indirect attention being given to all tokens. See Figure 2 for an illustrative example.
Averaging multiple attention matrices over different attention heads is a common approach to process or visualize the information contained in attention layers of transformers. This is despite numerous studies indicating that some heads are not informative for large transformers (Voita et al., 2019; Chefer et al., 2021b), suggesting that simple averaging might dilute the signal under observation. In the DTMC framework, averaging yields a new process where, at each step, the transitioning probabilities are the mean of those from the original chains — a “mixture” of the chains with equal weighting. This operation can be reasonable if the chains are not very different. As an alternative, we propose a weighting scheme that considers the convergence rate of the chains in Section 4.4.
@@ 4.2 Multi-Bounce Attention
When the subject of interest is a specific token $i$, the core shortcoming with row- and column-select from a Markovian perspective is that they do not take into account higher order effects: As illustrated in Figure 2, the tokens that token $i$ attends to are also attending other tokens, indirectly affecting token $i$. While it is evident that the attention mechanism only implements direct influence between tokens, we show that considering indirect paths allows for a deeper understanding of attention flow within individual attention heads. For brevity, we will now extend the row-select operation using our framework, which can be trivially applied to column-select by normalizing the columns of $\mathbf{A}$ and transposing.
@@
with $\mathbf{v}^{T}_{\{i,n=0\}}$ being a one hot row vector. Such an iterative process is mathematically equivalent to the power method (Equation 1). In other words, if we wish to take into account the $n^{\text{th}}$ order attention bounce, we can simply apply this formula $n$ times. For a single bounce, $\mathbf{v}^{T}_{\{i,n=1\}}$ captures incoming attention for tokens that $i$ attends to directly (row-select) or outgoing attention for tokens that attend $i$ directly (column-select). For any $\mathbf{v}^{T}_{\{i,n>1\}}$, higher order bounces are considered. See Figure 1 for an illustration.
@@ 4.3 TokenRank
In many cases, a more complete understanding of token importance is required. Observe that applying multi-bounce attention (Section 4.2) will converge into a stationary vector for any initial state if specific conditions are met (Section 3.1). In the context of attention matrices, we name this vector TokenRank, inspired by Google’s webpage stationary vector PageRank (Page et al., 1999). We argue that TokenRank is a better tool to measure the global importance of tokens in a single attention head, compared to more localized and naïve attempts to probe the attention matrix (e.g. by selecting a specific column, or summing up each of the columns).
One caveat is that attention weight matrices might still be reducible, or have cycles of fixed length. This means they do not necessarily hold the properties that guarantee the existence and uniqueness of a steady-state vector. A sufficient way to guarantee this, is to adjust $\mathbf{A}$ using the PageRank formulation (Equation 2) prior to iterating.
@@
We show that TokenRank extracts a cleaner signal of the attention (Section 5.2), can be used to improve downstream tasks (Section 5.3), and is better for determining the importance of tokens in the sequence compared to other approaches (Section 5.4).
@@ 4.4 $\lambda_{2}$ Weighting
The mixing time of a DTMC into its steady state is mathematically tied with its 2nd largest eigenvalue $\lambda_{2}$ (Langville and Meyer, 2004). Since DTMCs with randomly sampled transition probabilities tend to have extremely small $\lambda_{2}$ with little to no metastable states (see supplementary), we hypothesize that larger values are statistically correlated with the amount of important information captured by the attention matrix. To this end, instead of simply averaging attention matrices over the head dimension (Section 4.1), we instead propose to perform weighted averaging over the heads using their $\lambda_{2}$ values. We show this weighting scheme results in better downstream task performance (see Section 5.1 and supplementary material).
@@
Figure 3: ImageNet segmentation. Considering higher order attention effects improves results. We visualize the raw attention output (colored) and the binary segmentation masks. We present more qualitative comparisons in the supplementary.
@@ 5 Experiments
This section provides extensive evidence that our framework and the extended operations over the attention matrices are useful for a variety of downstream tasks. We start with improving zero-shot segmentation over a standard benchmark using multi-bounce attention in Section 5.1, followed by showing that TokenRank extracts a more informative signal than existing baselines in Section 5.2, useful for visualization purposes. Then, we demonstrate that applying TokenRank to an existing technique improves unconditional generation of images in Section 5.3. This is supplemented with a token masking experiment Section 5.4, where we deliberately mask tokens according to their TokenRank importance and measure downstream accuracy for image classification.
@@ 5.1 ImageNet Zero-Shot Segmentation
Table 1: ImageNet segmentation. Our method yields state-of-the-art results.
@@
A straightforward way to evaluate the usefulness of taking into account higher-order attention effects is to measure the performance on a downstream task such as zero-shot segmentation. We evaluate on the ImageNet Segmentation benchmark (Chefer et al., 2021b) consisting of 4276 images with 445 categories and we employ FLUX-Schnell (Labs, 2024), a pretrained transformer-based generative model, to denoise the images and extract the attention maps. We then average the maps over the heads using $\lambda_{2}$ weighting (Section 4.4) and compute the multi-bounce attention when considering outgoing links with $n=2$ (Section 4.2). We refer to the supplementary for further implementation details and ablations.
We report the results for the current state-of-the-art method (Helbling et al., 2025) and additionally provide the row- and column-select operations as baselines using their implementation. The results are reported in Table 1 and example qualitative segmentation masks are presented in Figure 13, indicating that our method is on par and arguably better than the previous state-of-the-art, as the threshold-agnostic metric mAP shows a large gap in performance. We hypothesize that this is because semantically similar regions form sets of metastable states, and iterating consolidates them, which results in cleaner segmentation maps. See supplementary for an illustrative example.
5.2 Linear Probing of TokenRank
Accuracy DINOv1 DINOv2 CLIP
Rand. Token 66.78 67.60 59.87
Center Token 65.81 79.01 69.96
Column Sum 75.52 91.34 75.80
TokenRank 77.50 93.27 76.13
CLS Token 79.92 95.11 74.90
@@ 5.2 Linear Probing of TokenRank
Figure 4: Global incoming attention. Visualizations are computed after averaging over heads for four different layers of DINOv1. While the center token only attends to the local neighborhood for earlier layers, column sum results in noisy attention visualizations. In contrast, TokenRank captures global incoming attention on par with the CLS token that was explicitly trained to capture global attention for DINOv1. We show per-head visualizations in the supplementary.
Previous typical approaches for visualizing global incoming attention consider the attention with respect to a Center Token, that is, row-select for that token (Hatamizadeh et al., 2024), or compute a per-Column Sum (Hong et al., 2023). These methods share the common characteristic that they can each be expressed as a single power-method iteration (Section 4.1). In other words, they only take into account direct attention effects. On the other hand, TokenRank considers indirect attention paths to evaluate where attention is flowing into. This results in attention maps that are sharper and less noisy than with column sum and more complete than using a single token (Figure 4).
To quantitatively evaluate that TokenRank produces sharper maps and captures the global information contained in the attention matrices, we train a linear classifier on top of all proposed attention visualizations, inspired by a previous similar analysis (Liu et al., 2024). For this purpose, we extract attention visualizations for all layers, heads, and images for the ImageNet (Deng et al., 2009) subset Imagenette (Howard, 2019) for three foundation models without considering the CLS token. The results in Table 2 show that TokenRank visualizations result in a cleaner signal for image classification than previous approaches that only consider one bounce of attention. It is on par with row-selecting the CLS token, which is an upper bound because it was specifically trained to capture class concepts for the DINO model family. See supplementary for further implementation details.

### v2

@@ Abstract
We introduce a new interpretation of the attention matrix as a discrete-time Markov chain. Our interpretation sheds light on common operations involving attention scores such as selection, summation, and averaging in a unified framework. It further extends them by considering indirect attention, propagated through the Markov chain, as opposed to previous studies that only model immediate effects. Our key observation is that tokens linked to semantically similar regions form metastable states, i.e., regions where attention tends to concentrate, while noisy attention scores dissipate. Metastable states and their prevalence can be easily computed through simple matrix multiplication and eigenanalysis, respectively. Using these lightweight tools, we demonstrate state-of-the-art zero-shot segmentation. Lastly, we define TokenRank—the steady state vector of the Markov chain, which measures global token importance. We show that TokenRank enhances unconditional image generation, improving both quality (IS) and diversity (FID), and can also be incorporated into existing segmentation techniques to improve their performance over existing benchmarks. We believe our framework offers a fresh view of how tokens are being attended in modern visual transformers.
@@ 1 Introduction
We introduce the interpretation of individual attention matrices as discrete-time Markov chains (DTMC) (Figure˜1), where the non-negative attention weights indicate transition probabilities between states that correspond to tokens. This framework, dubbed as Attention Chains, allows analyzing token importance through multiple chain transitions, even though the attention operation models only direct relationships. This is analogous to Google’s seminal paper on DTMC, PageRank (Page et al., 1999), stating that simple counting of incoming hyperlinks does not generally correspond to page importance. Similarly, simply using direct attention to measure global relevance of a token under performs compared to propagating it through the chain, as we show in various experiments.
@@
Based on this insight, we propose a simple yet powerful operation—considering several transitions, or bounces, of the Markov Chain. This consolidates metastable states and helps filter out noise in the attention signal. As we demonstrate, using multiple bounces yields state-of-the-art results for zero-shot segmentation on a commonly used benchmark. Metastable states can also be identified via eigenanalysis, and their presence relates to higher 2nd largest eigenvalues. We use this characteristic to weigh attention heads and improve image segmentation.
@@ 2 Related Work
Due the importance of the attention operation in transformers, studies have proposed modeling them using various mathematical frameworks. For example, Ramsauer et al. (2020) relate attention to classic Hopfield networks, Tsai et al. (2019) frame attention as a kernel smoother, Schlag et al. (2021) find its relation to classic fast weight controllers, or Raghu et al. (2021) compare vision transformers to CNNs. El et al. (2025) discuss the mathematical equivalence between message passing in graph neural networks and the self-attention operation in transformers. Ildiz et al. (2024) consider modeling a 1-layer transformer as a Markov chain. In contrast, we focus on large pretrained transformers and model each individual self-attention matrix as a Markov chain. While their approach allows explaining some typical behaviors of transformers, we explicitly show how our framework allows for the interpretation of existing, pretrained large transformers and for better performance in various downstream tasks. Secondly, we leverage more of the machinery of the Markov chain theory, including multiple transitions and steady state analysis, which was not considered previously.
@@ 3.1 Discrete-Time Markov Chains (DTMC)
where $\mathbf{v}^{T}_{n=0}$ is an initial state summing up to one. The computation is usually terminated once the norm falls below a specific threshold $\mathbf{v}^{T}_{n+1}-\mathbf{v}^{T}_{n}||^{2}_{2}<\tau$ or after some fixed number of iterations.
@@ 3.2 PageRank Algorithm
where $0<\alpha<1$ is a hyper parameter controlling the transitioning probability of “teleportation” into a random state, and $\mathbf{e}$ is a column vector of all ones. Applying the power method on $\mathbf{P}^{\prime}$ results in a unique vector $\vec{\mathbf{\pi}}$, the global PageRank vector, inducing the order $i\preceq_{\vec{\mathbf{\pi}}}j\quad\Longleftrightarrow\quad\vec{\mathbf{\pi}}_{i}\leq\vec{\mathbf{\pi}}_{j}$.
Figure 2: Illustration of higher order effects. Left: Attention matrix $A$ with sequence length 5. Middle: A DTMC with transition probabilities defined by matrix AA, where only strong connections are shown. Right (One-Hot): To evaluate where state-4 attends to, we can iterate using the power method once starting from a one-hot vector ($n=0$), which results in the row-select operation ($n=1$). However, this first-order approximation is insufficient since state-0 mostly transitions to state-3 and, therefore, state-4 indirectly attends state-3. This becomes evident as we iterate further ($n=2$). Right (Uniform): To compute a global token ranking, we can iterate starting from a uniform state ($n=0$), resulting in a per-column sum operation ($n=1$). This indicates state-0 as most important because many states have a high probability of transitioning into state-0. However, state-0 maps to state-3 with high probability, and state-3 maps to state-4 with high probability. Therefore, the importance of state-4 should be elevated. When considering the second bounce ($n=2$), more probability mass is directed into state-3, and with a sufficient number of iterations the steady state ($v^{T}_{ss}$) ranks state-4 as the most important state globally, which aligns with the intuition above.
@@ 4 Attention as a Discrete-Time Markov Chain
where $h$ is the number of heads, $s_{1,2}$ are the token sequence lengths, and $d_{h}$ is a sub-space of the full embedding dimension $d$. Consequently, we interpret the attention matrix as a DTMC, going beyond the conventional attention mechanism in transformers, which only models direct interaction between queries and keys. Our discussion is limited to attention blocks of equal sequence size (e.g. self-attention, or hybrid attention blocks), unless otherwise stated. We start by formalizing some common operations performed on attention matrices using the DTMC formalism (Section˜4.1). While the previously proposed attention operations can be formulated with a single bounce of the Markov chain, we go beyond this and iterate through the Markov chain. Specifically, we describe three new operations enabled by this framework: multi-bounce attention (Section˜4.2); where higher-order attention effects are taken into account, TokenRank (Section˜4.3); the token-space equivalent of PageRank, and $\lambda_{2}$ weighting; an improved attention head weighting scheme that is based on the second largest eigenvalue of the chain.
@@ 4.1 Interpretation of Existing Attention Operations
Thus, our formulation allows us to reinterpret the selection operation as the first bounce transition from a one-hot vector, a view that we will further motivate in Section˜4.2.
Summing each column of $\mathbf{A}$ results in a row vector $\mathbf{v_{s}}^{T}$ that aggregates attention given to each token from all other tokens. In our formulation, $\mathbf{v_{s}}^{T}$ can be obtained by performing a single state transition: $\mathbf{v}^{T}_{s}=\frac{1}{n}\mathbf{e}^{T}\mathbf{A}$, where $\mathbf{e}^{T}$ is a row vector of ones. This operation can be a useful first-order approximation in ranking tokens’ importance globally, as it equates to starting from an equilibrium state and transitioning through the DTMC process once. However, we show this straight-forward approach under-performs compared to the global TokenRank vector (Section˜4.3) that takes into account indirect attention being given to all tokens. See Figure˜2 for an illustrative example.
Averaging multiple attention matrices over different attention heads is a common approach to process or visualize the information contained in attention layers of transformers. This is despite numerous studies indicating that some heads are not informative for large transformers (Voita et al., 2019; Chefer et al., 2021b), suggesting that simple averaging might dilute the signal under observation. In the DTMC framework, averaging yields a new process where, at each step, the transitioning probabilities are the mean of those from the original chains — a “mixture” of the chains with equal weighting. This operation can be reasonable if the chains are not very different. As an alternative, we propose a weighting scheme that considers the convergence rate of the chains in Section˜4.4.
@@ 4.2 Multi-Bounce Attention
When the subject of interest is a specific token $i$, the core shortcoming with row- and column-select from a Markovian perspective is that they do not take into account higher order effects: As illustrated in Figure˜2, the tokens that token $i$ attends to are also attending other tokens, indirectly affecting token $i$. While it is evident that the attention mechanism only implements direct influence between tokens, we show that considering indirect paths allows for a deeper understanding of attention flow within individual attention heads. For brevity, we will now extend the row-select operation using our framework, which can be analogously applied to column-select after normalizing the columns of $\mathbf{A}$ and transposing.
@@
with $\mathbf{v}^{T}_{\{i,n=0\}}$ being a one hot row vector. Such an iterative process is mathematically equivalent to the power method (Equation˜1). In other words, if we wish to take into account the $n^{\text{th}}$ order attention bounce, we can simply apply this formula $n$ times. For a single bounce, $\mathbf{v}^{T}_{\{i,n=1\}}$ captures incoming attention for tokens that $i$ attends to directly (row-select) or outgoing attention for tokens that attend $i$ directly (column-select). For any $\mathbf{v}^{T}_{\{i,n>1\}}$, higher order bounces are considered. See Figure˜1 for an illustration.
@@ 4.3 TokenRank
In many cases, a more complete understanding of token importance is required. Observe that applying multi-bounce attention (Section˜4.2) will converge into a stationary vector for any initial state if specific conditions are met (Section˜3.1). In the context of attention matrices, we name this vector TokenRank, inspired by Google’s webpage stationary vector PageRank (Page et al., 1999). We argue that TokenRank is a better tool to measure the global importance of tokens in a single attention head, compared to more localized and naïve attempts to probe the attention matrix (e.g. by selecting a specific column, or summing up each of the columns).
One caveat is that attention weight matrices might still be reducible, or have cycles of fixed length. This means they do not necessarily hold the properties that guarantee the existence and uniqueness of a steady-state vector. A sufficient way to guarantee this, is to adjust $\mathbf{A}$ using the PageRank formulation (Equation˜2) prior to iterating.
@@
We show that TokenRank extracts a cleaner signal of the attention (Section˜5.2), can be used to improve downstream tasks (Section˜5.3), and is better for determining the importance of tokens in the sequence compared to other approaches (Section˜5.5).
@@ 4.4 $\lambda_{2}$ Weighting
The mixing time of a DTMC into its steady state is mathematically tied with its 2nd largest eigenvalue $\lambda_{2}$ (Langville and Meyer, 2004). Intuitively, larger $\lambda_{2}$ values correspond to more stable metastable states, indicating that more important information is captured by the attention matrix. Indeed, DTMCs with randomly sampled transition probabilities tend to have very small $\lambda_{2}$ (see Appendix˜F).
To this end, instead of simply averaging attention matrices over the head dimension (Section˜4.1), we instead propose to perform weighted averaging over the heads using their $\lambda_{2}$ values. We show this weighting scheme results in better downstream task performance (see Section˜5.1 and Section˜E.2).
@@
Figure 3: ImageNet segmentation. Considering higher order attention effects improves results. We visualize the raw attention output (colored) and the binary segmentation masks. We present more qualitative comparisons in Section˜C.2 .
@@ 5 Experiments
This section provides extensive evidence that our framework and the extended operations over the attention matrices are useful for a variety of downstream tasks. We start with improving zero-shot segmentation over a standard benchmark using multi-bounce attention (Section˜5.1), followed by showing that TokenRank extracts a more informative signal than existing baselines (Section˜5.2), useful for visualization purposes. Then, we demonstrate that applying TokenRank to existing techniques improves unconditional generation of images (Section˜5.3) and segmentation (Section˜5.4). This is supplemented with a token masking experiment (Section˜5.5), where we deliberately mask tokens according to their TokenRank importance and measure downstream accuracy for image classification.
@@ 5.1 ImageNet Zero-Shot Segmentation
Table 1: ImageNet segmentation. Our method yields state-of-the-art results. Error bars and further results are reported in Appendix˜C.
@@
A straightforward way to evaluate the usefulness of taking into account higher-order attention effects is to measure the performance on a downstream task such as zero-shot segmentation. We evaluate on the ImageNet Segmentation benchmark (Chefer et al., 2021b) consisting of 4276 images with 445 categories and we employ FLUX-Schnell (Labs, 2024), a pretrained transformer-based generative model, to denoise the images and extract the attention maps. We then average the maps over the heads using $\lambda_{2}$ weighting (Section˜4.4) and compute the multi-bounce attention when considering outgoing links with $n=2$ (Section˜4.2). We refer to the Appendix˜A for further implementation details and to Appendix˜C for further results.
We report the results for the current state-of-the-art method (Helbling et al., 2025) and additionally provide the row- and column-select operations as baselines using their implementation. The results are reported in Table˜1 and example qualitative segmentation masks are presented in Figure˜3, indicating that our method is better than the previous state-of-the-art, as the threshold-agnostic metric mAP shows a large gap in performance. We hypothesize that this is because semantically similar regions form sets of metastable states, and iterating consolidates them, which results in cleaner segmentation maps. Our proposed $\lambda_{2}$ weighting further improves the segmentation performance, as we show in statistical tests in Section˜E.1.
@@
Figure 4: Global incoming attention. Visualizations are computed after averaging over heads for four different layers of DINOv1. While the center token only attends to the local neighborhood for earlier layers, column sum results in noisy attention visualizations. In contrast, TokenRank captures global incoming attention on par with the CLS token that was explicitly trained to capture global attention for DINOv1. We show per-head visualizations in Appendix˜H.
5.2 TokenRank for Visualizing Attention Maps
Previous typical approaches for visualizing global incoming attention consider the attention with respect to a Center Token, that is, row-select for that token (Hatamizadeh et al., 2024), or compute a per-Column Sum (Hong et al., 2023). These methods share the common characteristic that they can each be expressed as a single power-method iteration (Section˜4.1). In other words, they only take into account direct attention effects. On the other hand, TokenRank considers indirect attention paths to evaluate where attention is flowing into. This results in attention maps that are sharper and less noisy than with column sum and more complete than using a single token (Figure˜4). We provide a quantitative comparison through a linear probing experiment Appendix˜B.
5.3 Improving Self-Attention Guidance (SAG)
To evaluate TokenRank on a downstream task, we integrate it in SAG (Hong et al., 2023), an approach that improves the fidelity and the quality of unconditional image generation with denoising diffusion models. SAG uses the self-attention weight matrices to form a mask that indicates important spatial tokens. It then adversarially blurs the masked regions and drives the denoising process away from it, creating better images as a result. To find out where important features exist, Hong et al. (2023) average over the head dimension for a specific attention layer, followed by a summing each of the columns (Section˜4.1). We compare the quality and diversity of generated images when using the TokenRank instead of their proposed Column-sum strategy. We compute the IS (Salimans et al., 2016), FID (Heusel et al., 2017), and KID (Bińkowski et al., 2018) metrics over 50K generated images. The selected metrics offer complementary information in terms of quality, diversity, and resemblance to ground-truth distribution. Results are reported in Table˜2 and Figure˜5. We refer to the supplementary material for further details and results. Using TokenRank improves the generated images significantly in both quality and diversity compared to the original SAG, showcasing TokenRank’s ability to rank globally important tokens in the sequence, and improve a downstream generative task.